An office supply company makes notebooks. The MRP for notebooks is given below.

notebooks

week1

2

3

4

5

6

gross req.

25

25

25

25

25

25

sch.recpt.

pr.av.bal.

30

5

0

0

0

0

0

pl.ord.recpt.

20

25

25

25

25

pl.ord.rels.

20

25

25

25

25

(Order quantity = L4L. Lead time = 1. Safety stock = 0.)

Verify that the MRP table is correct.  

Notebook assembly is carried out in workcenter 100, and needs a set up time of 1.8 hours, and a run time of 9 minutes per book (in the week when the order is received).

In week 2, how many minutes are needed at the workcenter 100?
Answer #1

MRP table is correct

pr. av. bal. (t+1) = pr. av. bal.(t) - Gross req. (t+1) + pl ord. recpt. (t+1)

In week 2, order is placed of 25 units

Number of minutes required = Set up time in minutes + Run time per book*Number of books ordered = 1.8*60 + 9*25 = 108 + 225 = 333 minutes